### v1

# Frequency Regulation for Exposure Bias Mitigation in Diffusion Models

Abstract.
Diffusion models exhibit impressive generative capabilities but are significantly impacted by exposure bias. In this paper, we make a key observation: the energy of the predicted noisy images decreases during the diffusion process. Building on this, we identify two important findings: 1) The reduction in energy follows distinct patterns in the low-frequency and high-frequency subbands; 2) This energy reduction results in amplitude variations between the network-reconstructed clean data and the real clean data. Based on the first finding, we introduce a frequency-domain regulation mechanism utilizing wavelet transforms, which separately adjusts the low- and high-frequency subbands. Leveraging the second insight, we provide a more accurate analysis of exposure bias in the two subbands. Our method is training-free and plug-and-play, significantly improving the generative quality of various diffusion models and providing a robust solution to exposure bias across different model architectures. The source code is available at https://github.com/kunzhan/wpp.
Diffusion model, exposure bias, wavelet transform
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06submissionid: 4742ccs: Generative model diffusion modelccs: Computing methodologies Computer vision problems
1. Introduction
In recent years, Diffusion Probabilistic Models (DPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) have made remarkable progress in image generation (Dhariwal and Nichol, 2021; Rombach et al., 2022). ADM (Dhariwal and Nichol, 2021) made the generation performance of DPMs surpass GAN (Goodfellow et al., 2014) by introducing classifier guidance. EDM (Karras et al., 2022) Significantly promotes development by clarifying the design space of DPMs. Additionally, EA-DDPM (Bao et al., 2022a), PFGM++ (Xu et al., 2023), and AMED (Zhou et al., 2024) also promote the improvement of DPMs from different aspects. However, all of these models still suffer from exposure bias. The exposure bias (Ning et al., 2023) is defined as the mismatch between training and inference in DPMs.
(a) The low-frequency energy.
(b) The high-frequency energy.
Figure 1. The subband energy of the noisy images in training and sampling, with the denoising direction being from right to left.
Exposure bias remains a significant challenge in diffusion models, and a variety of methods have been proposed to mitigate its impact. These approaches include diverse training strategies (Ning et al., 2023; Li and van der Schaar, 2024; Ren et al., 2024) and sampling techniques (Li et al., 2024; Zhang et al., 2025; YAO et al., 2025). However, the majority of these methods predominantly operate in the spatial domain, neglecting the potential of addressing exposure bias through exploration in the frequency domain. As illustrated in Fig. 1, the energy of the predicted noisy images consistently decreases during the diffusion process, with distinct patterns of energy reduction observed across different frequency subbands. By applying wavelet transform(Graps, 1995), we analyze the evolution of exposure bias within the frequency domain and uncover two key findings: 1) The energy reduction in the low-frequency subband persists throughout the sampling process, and 2) High-frequency energy is primarily diminished towards the later stages of the diffusion process.
Furthermore, previous research has largely overlooked the mathematical formulation of exposure bias (Ning et al., 2024; Li et al., 2024), relying on strong assumptions to approximate its impact. Through detailed analytical modeling, we validate our observations and derive the precise mathematical expression for exposure bias, providing a clearer understanding of its underlying mechanisms in two subbands.
Through discrete wavelet transform (DWT), the predicted noisy image is decomposed into high-frequency and low-frequency subbands. The low-frequency subband is then amplified throughout the entire sampling process using a low-frequency adjustment weight, while the high-frequency subbands are amplified towards the end of the sampling process using a high-frequency weight. Finally, the adjusted subbands are combined and reconstructed into the predicted noisy images in the spatial space via inverse DWT (iDWT).
However, we observe that amplifying the low-frequency component towards the end of the sampling process may conflict with the denoising principles of DPMs, potentially introducing undesirable effects. Specifically, DPM aims to finely restore image details in the later stages of sampling. As shown in Fig. 2, high-frequency energy begins to change significantly only at the final stages. If a relatively large weight is assigned to the low-frequency component during this phase, it hinders the refinement process. To address this issue, we explore several dynamic weighting strategies, where the influence of the low-frequency component decreases over time, while the high-frequency component is gradually amplified as sampling progresses.
[…]
In summary, our contributions are:
  • We are, to the best of our knowledge, the first to analyze and solve the exposure bias of DPMs from the perspective of the frequency domain. Through analytical modeling and extensive experiments, we derive for the first time a rigorous mathematical form of exposure bias.
  • We propose a frequency-domain regulation mechanism based on wavelet transform to mitigate exposure bias, which is train-free and plug-and-play, improving the generation quality of various DPMs with negligible computational cost.
  • Our method can achieve high-quality performance metrics with negligible computational cost, outperforming all current improved models for exposure bias.
The rest of this paper is organized as follows. In Section 2, we review related work on exposure bias in diffusion models and methods to mitigate it. Section 3 presents the preliminaries, including background on diffusion models, exposure bias, and wavelet transform. In Section 4, we introduce our proposed method for mitigating exposure bias using frequency-domain regulation and dynamic weighting. Section 5 details the experiments conducted to evaluate our approach, followed by a discussion of the results. Finally, Section 6 concludes the paper and suggests potential directions for future work.
2. Related Work
This section reviews prior research on exposure bias in diffusion models, highlighting existing methods and approaches that aim to mitigate this issue. We also discuss relevant work in frequency-domain analysis, wavelet transforms, and techniques for improving generative models, positioning our work within the context of these studies.
The theoretical foundation of DPMs was established by DPM (Sohl-Dickstein et al., 2015), with substantial advancements later achieved by DDPM (Ho et al., 2020). By introducing classifier guidance, ADM (Dhariwal and Nichol, 2021) made the generation performance of DPMs surpass GAN (Goodfellow et al., 2014) for the first time. SDE (Song et al., 2021b) proposes a unified DPM framework by means of stochastic differential equations, while DDIM (Song et al., 2021a) accelerates sampling by skipping steps. Analytic-DPM (Bao et al., 2022b) theoretically derived the optimal variance in the reverse process, while EDM (Karras et al., 2022) presented a design space to separate the concrete design choices in DPMs. In addition, ODE-based DPMs (Lu et al., 2022; Zhou et al., 2024; Zhao et al., 2024; Zhang and Chen, 2023; Dockhorn et al., 2022), consistency models (Song et al., 2023; Song and Dhariwal, 2024; Lu and Song, 2025) and distillation techniques (Salimans and Ho, 2022; Liu et al., 2023; Meng et al., 2023; Luhman and Luhman, 2021; Zheng et al., 2023) are widely developed, which significantly speeds up sampling.
Wavelet transform (Graps, 1995; Mallat, 1989) has been widely applied in the field of generative adversarial networks (Gal et al., 2021; Yang et al., 2022; Wang et al., 2022; Zhang et al., 2022), improving the quality of various generation tasks. In recent years, some works have started to combine wavelet transform with diffusion models. WACM  (Li et al., 2022) utilizes the wavelet spectrum to assist image colorization. WSGM (Guth et al., 2022) factorizes the data distribution into the product of conditional probabilities of wavelet coefficients at various scales to accelerate the generation process. WaveDiff (Phung et al., 2023) processes images in the frequency domain to further enhance the efficiency and quality of generation.
The exposure bias in DPMs was first systematically analyzed by ADM-IP (Ning et al., 2023), which simulated exposure bias by re-perturbing the training sample. Subsequently, EP-DDPM (Li and van der Schaar, 2024) regard the cumulative error as a regularization term to mitigate exposure bias, while MDSS (Ren et al., 2024) employed the multi-step denoising scheduled sampling strategy to mitigate this issue. It should be noted these three methods require retraining of the model. TS-DPM (Li et al., 2024) put forward a time-shift sampler and ADM-ES applied the noise scaling technique, which both can mitigate exposure bias without training. Additionally, AE-DPM (Zhang et al., 2025) alleviates exposure bias through prompt learning, whereas MCDO (YAO et al., 2025) mitigates it by means of manifold constraint.
Unfortunately, all the previous work on exposure bias was grounded in expansion within the spatial space. In contrast, we analyze this issue from the point of view of the frequency domain and address it using frequency-domain approaches.
3. Preliminaries
In this section, we provide the necessary background for understanding our approach. We introduce diffusion models, the concept of exposure bias, and the fundamentals of wavelet transform. This section lays the foundation for the detailed methodology presented in the next section. We first introduce the diffusion model (DPM) (Ho et al., 2020), followed by the exposure bias (Ning et al., 2023) and the wavelet transform.
3.1. Diffusion model
DPM is divided into forward noising and reverse denoising process, both of which are defined as Markov processes. For a original data distribution $q({\bm{x}}_{0})$ and a noise schedule $\beta_{t}$, the forward process is
(1) $q\left({\bm{x}}_{1:T}|{\bm{x}}_{0}\right)=\prod_{t=1}^{T}q\left({\bm{x}}_{t}|{\bm{x}}_{t-1}\right),$
where $q\left({\bm{x}}_{t}|{\bm{x}}_{t-1}\right)=\mathcal{N}\left({\bm{x}}_{t};\sqrt{1-\beta_{t}}{\bm{x}}_{t-1},\beta_{t}{\bm{I}}\right)$. Specifically, by leveraging the properties of the Gaussian distribution, the perturbed distribution can be rewritten as the conditional distribution $q\left({\bm{x}}_{t}|{\bm{x}}_{0}\right)$, which is expressed by the formula
(2) ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{t},$
where ${\alpha}_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and ${\bm{\epsilon}}_{t}\sim\mathcal{N}(\bm{0},{\bm{I}})$. Based on Bayes’ theorem, the posterior conditional probability is obtained, which is
(3) $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})=\mathcal{N}(\tilde{\mu}_{t}({\bm{x}}_{t},{\bm{x}}_{0}),\tilde{\beta}_{t}{\bm{I}}),$
where $\tilde{\mu}_{t}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}{\bm{x}}_{0}+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}{\bm{x}}_{t}$ and $\tilde{\beta}_{t}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}$. We apply a neural network $p_{{{\bm{\theta}}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}({\bm{x}}_{t-1};\mu_{{\bm{\theta}}}({\bm{x}}_{t},t),\sigma_{t}{\bm{I}})$ to approximate $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$, and naturally we want to minimize $D_{\rm KL}(q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||p_{{\bm{\theta}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})))$. Through a series of reparameterization, we obtain
(4) $\mu_{{\bm{\theta}}}({\bm{x}}_{t},t)$ $=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}{\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}{\bm{x}}_{t}$
$=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)\right),$
where ${\bm{\epsilon}}_{\bm{\theta}}(\cdot)$ is a noise prediction network and ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ represent a reconstruction model which predicts ${\bm{x}}_{0}$ given ${\bm{x}}_{t}$:
(5) ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)=\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$
So the final loss function is given by
(6) $\mathcal{L}_{\rm simple}=\mathbb{E}_{t,{\bm{x}}_{0},{\bm{\epsilon}}_{t}\sim\mathcal{N}(\mathbf{0},{\bm{I}})}[\|{\bm{\epsilon}}_{{\bm{\theta}}}({\bm{x}}_{t},t)-{\bm{\epsilon}}_{t}\|_{2}^{2}].$
Once ${\bm{\epsilon}}_{{\bm{\theta}}}(\cdot)$ has converged, we can perform reverse sampling starting from a standard Gaussian distribution by using $p_{{{\bm{\theta}}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})$.
3.2. Exposure bias
Specifically, for a pre-trained noise prediction network ${\bm{\epsilon}}_{\bm{\theta}}(\cdot)$, the network input comes from the corrupted noisy image ${\bm{x}}_{t}$ based on the ground truth during training, while during sampling the network input comes from the predicted noisy image $\hat{{\bm{x}}}_{t}$. Due to the prediction error of the neural network and the discretization error of the numerical solver, the sampling trajectory of the DPM tends to deviate from the ideal path, which leads to the deviation between the predicted noisy image $\hat{{\bm{x}}}_{t}$ and the corrupted sample ${\bm{x}}_{t}$. This exposure bias between noisy images leads to a bias in the network output ${\bm{\epsilon}}_{\bm{\theta}}\left(\hat{{\bm{x}}}_{t},t\right)$ and ${\bm{\epsilon}}_{\bm{\theta}}\left({\bm{x}}_{t},t\right)$, which exacerbates the exposure bias in the next time step. Therefore, the exposure bias gradually accumulates with the sampling iteration, and ultimately affect the generation quality as shown in Fig. 3.
Figure 3. The corrupted noisy image ${\bm{x}}_{t}$ is generated from the clean image ${\bm{x}}_{0}$ during training through the diffusion process. However, during inference, due to discretization errors in the SDE or ODE solvers, as well as the inherent prediction errors in the network, the predicted noisy image $\hat{{\bm{x}}}_{t}$ deviates from the ideal trajectory. This deviation leads to a discrepancy between ${\bm{x}}_{t}$ and $\hat{{\bm{x}}}_{t}$, causing a consistent bias in the predicted noisy image. As a result, ${\bm{x}}_{t}$ and $\hat{{\bm{x}}}_{t}$ are always biased, with the extent of this bias being influenced by both the solver’s discretization and the network’s prediction accuracy.
The exposure bias is the mismatch between forward and reverse in DPMs. For a frozen noise prediction network ${\bm{\epsilon}}_{\bm{\theta}}(\cdot)$, the input to the network is an ideal corrupted noisy image ${\bm{x}}_{t}$ based on the ground truth $x_{0}$ by Eq. (2) in the forward process. However, during the reverse process, the input to network is an actual prediction noisy image $\hat{{\bm{x}}}_{t}$, which depends on the prediction noisy image of previous timestep. Due to the prediction error of the neural network and the discretization error of the numerical solver, the sampling trajectory tends to deviate from the ideal trajectory, which leads to the difference between the predicted noisy image $\hat{{\bm{x}}}_{t}$ and the corrupted noisy image ${\bm{x}}_{t}$. Thus, due to exposure bias, the actual sampling formula should be
(7) $\hat{{\bm{x}}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\hat{{\bm{x}}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)\right)+\sigma_{t}{\bm{z}}.$
Specially, Both $\hat{{\bm{x}}}_{T}$ and ${\bm{x}}_{T}$ obey the standard Gaussian distribution, so there is no bias. However, due to the prediction error between ${\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{T},T)$ and the ideal noise, the sampling operation in Eq. (7) leads to the deviation between $\hat{{\bm{x}}}_{T-1}$ and ${\bm{x}}_{T-1}$, which further exacerbates the bias between ${\bm{\epsilon}}_{\bm{\theta}}(\hat{{\bm{x}}}_{T-1},T)$ and ${\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{T-1},T)$. Thus, as sampling continues, the bias of the predicted noisy image and the bias of the network influences each other and gradually accumulate, causing the exposure bias between $\hat{{\bm{x}}}_{t}$ and ${\bm{x}}_{t}$.
3.3. Wavelet transform
We introduce DWT which decomposes an image ${\bm{x}}\in\mathbb{R}^{H\times W}$ into four wavelet subbands, namely ${\bm{x}}^{ll},{\bm{x}}^{lh},{\bm{x}}^{hl},{\bm{x}}^{hh}$. Conversely, the wavelet subbands can be reconstructed into an image by using the iDWT. The two processes are expressed as:
(8) $\begin{cases}\{{\bm{x}}^{f}|f=ll,lh,hl,hh\}=\textrm{DWT}({\bm{x}})\\ {\bm{x}}=\textrm{iDWT}({\bm{x}}^{f}|f=ll,lh,hl,hh)\end{cases}$
where the size of all four subbands is $\mathbb{R}^{H/2\times W/2}$. ${\bm{x}}^{ll}$ represents the low-frequency component of the image ${\bm{x}}$, reflecting the basic structure of the image, such as the human face shape and the shape of birds. ${\bm{x}}^{lh},{\bm{x}}^{hl}$, and ${\bm{x}}^{hh}$ represent the high-frequency components in different directions of the image, reflecting the detailed information of the image, such as human wrinkles and bird feathers. For simplicity, we choose the classic Haar wavelet in experiments.
4. methods
In this section, we begin by analyzing exposure bias from a frequency-domain perspective using wavelet transform. Our key observation reveals that exposure bias leads to the gradual reduce of low-frequency energy throughout the sampling process, and the reduction of high-frequency energy in the final stages of sampling. To validate this observation, we conduct analytical modeling, confirming the distinct patterns of energy reduction in both low- and high-frequency subbands. Based on these findings, we propose a simple yet effective frequency-energy regulation mechanism designed to mitigate exposure bias. This mechanism operates by adjusting the frequency energy distribution at each timestep of the sampling process. Furthermore, we introduce a novel dynamic weighting scheme, which enables more accurate scaling of frequency energy, allowing the low-frequency components to be emphasized in the early stages of sampling, while the high-frequency components are given more weight towards the end. Together, these findings and strategies form a robust approach to addressing exposure bias, improving the generative performance of diffusion models.
4.1. Energy reduction in frequency space
(a) The $ll$ subband of ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$
(b) The $lh$ subband of ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$
Figure 4. The subband energy of ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$ in training and sampling.
Now, we officially analyze exposure bias in the frequency domain.
Firstly, we need to simulate exposure bias accurately. During sampling, it is easy to obtain the predicted noisy image $\hat{{\bm{x}}}_{t}$, but acquiring the ideal perturbed noisy image ${\bm{x}}_{t}$ corresponding to $\hat{{\bm{x}}}_{t}$ is difficult. To solve this problem, we rely on the original data and the deterministic solver to simulate the exposure bias. Specifically, for the given original data ${\bm{x}}_{0}$, we first perform forward perturbation on ${\bm{x}}_{0}$ to obtain a series of perturbed noisy images ${\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{s+1}$. Then, we take ${\bm{x}}_{s+1}$ as the actual sampling starting point, match the corresponding time step $s+1$ and coefficients, and carry out the DDIM sampling
(9) $\hat{{\bm{x}}}_{s-1}=\sqrt{\bar{\alpha}_{s-1}}\left(\frac{x_{s}-\sqrt{\bar{\alpha}_{s}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{s},s)}{\sqrt{\bar{\alpha}_{s}}}\right)+\sqrt{1-\bar{\alpha}_{s-1}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{s},s).$
As a result, we obtain a series of noisy images $\hat{{\bm{x}}}_{s-1},\ldots,\hat{{\bm{x}}}_{2},\hat{{\bm{x}}}_{1},\hat{{\bm{x}}}_{0}$. As long as we ensure ${\bm{x}}_{s}$ still maintain a certain level of low-frequency information of ${\bm{x}}_{0}$, we are able to make the reverse predicted noisy images $\hat{{\bm{x}}}_{t}$ similar to the forward perturbed noisy images ${\bm{x}}_{t}$.
Secondly, we perform Discrete Wavelet Transform (DWT) on the disturbed and predicted noisy images at different time steps to get their corresponding frequency components. Then, we calculate respectively the mean squared norm of the subbands of these two batches of noisy images at different time steps. Fig. 1(a) clearly shows that as the sampling progresses, the energy of the low-frequency subbands of $\hat{{\bm{x}}}_{t}$ gradually becomes lower than that of ${\bm{x}}_{t}$, which indicates there is a phenomenon of reducing of low-frequency energy during sampling. Fig. 1(b) reveals two important phenomena: (1) From the start of sampling to the end, the model does not prioritize the construction of high-frequency energy, resulting in a small deviation in high-frequency energy during this process. (2) At the end of sampling, the model significantly reduces high-frequency energy.
The above experiments indicate there is a reduction of low-frequency energy throughout the sampling process and a reduction of high-frequency energy at the end of sampling, which is the manifestation of exposure bias in the frequency domain.
Table 1. Mean and variance of $q({\bm{x}}_{t}|{\bm{x}}_{0})$ and $p_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t}|{\bm{x}}_{t+1})$
Mean Variance
$q({\bm{x}}_{t}|{\bm{x}}_{0})$ $\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}$ $(1-\bar{\alpha}_{t})\bm{I}$
$p_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t}|{\bm{x}}_{t+1})$ $\gamma_{t}\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}$ $\left(1-\bar{\alpha}_{t}+\left(\frac{\sqrt{\bar{\alpha}_{t}}\beta_{t+1}}{1-\bar{\alpha}_{t+1}}\phi_{t+1}\right)^{2}\right)\bm{I}$
4.2. Theoretical perspective
We hope to explain and prove the observations via analytical modeling. Previous work (Ning et al., 2024; Li et al., 2024) rely heavily on a strong assumption
(10) ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)={\bm{x}}_{0}+\phi_{t}{\bm{\epsilon}}_{t},$
where ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ represent the reconstruction model predicting ${\bm{x}}_{0}$ given ${\bm{x}}_{t}$, shown in Eq. (5), and $\phi_{t}$ is a scalar coefficient related to time steps. However our approach revises the assumption.
To explore the modeling rule of ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$, we used Eq. (5) on the basis of experiment 1 to obtain ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ and ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)$ corresponding to ${\bm{x}}_{t}$ and $\hat{{\bm{x}}}_{t}$, and performed the same wavelet decomposition and norm calculation with details for appendix. Fig. 4(a) and  4(b) clearly show that neither ${\bm{x}}_{t}$ nor $\hat{{\bm{x}}}_{t}$ can fully reconstruct ${\bm{x}}_{0}$ at any time step, and ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)$ reconstructed by the predicted noisy images is smaller than ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ reconstructed by the perturb noisy images regardless of the frequency component and time step. Based on this key observation, we make a more scientific assumption about $x_{0}$.
Assumption 1 ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)$ is assumed to be represented in terms of ${\bm{x}}_{0}$
(11) ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)=\gamma_{t}{\bm{x}}_{0}+\phi_{t}{\bm{\epsilon}}_{t},$
where ${\bm{\epsilon}}_{t}\sim\mathcal{N}(\bm{0},{\bm{I}})$, $0<\gamma_{t}\leqslant 1$, $\phi_{t}<M$, and $M$ is an upper bound. Based on this assumption, we re-parameterize the conditional probability in the sampling process:
(12) $\hat{{\bm{x}}}_{t-1}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}{\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}{\bm{x}}_{t}+\sqrt{\tilde{\beta_{t}}}{\bm{\epsilon}}_{1}.$
Through plugging Eq. (11) and Eq. (2) into Eq. (12), we obtain the analytical form of $\hat{{\bm{x}}}_{t}$:
(13) $\hat{{\bm{x}}}_{t-1}=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t-1}+\left(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t}\right)^{2}}{{\bm{\epsilon}}}_{2},$
where ${\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$. And the proof of  13 is provided in the appendix. Table 1 clearly shows that the predicted distribution always has a larger variance and a smaller mean than the perturbation distribution, indicating that prediction noisy images reduces energy about the target data ${\bm{x}}_{0}$ during sampling. Naturally, as the diffusion model focuses on recovering low-frequency energy in early stages, predictions mainly lose low-frequency components then; in late stages, when the model reconstructs high-frequency details, predictions primarily lose high-frequency components.
4.3. Frequency regulation
Based on previous key findings, we propose a simple and effective mechanism for frequency domain component regulation based on wavelet transform, as shown in Fig. 5. . Specifically, for the predicted noisy image $\hat{{\bm{x}}}_{t}$ at each step of the sampling process, we use the wavelet transform to decompose it into four distinct frequency components: $\hat{{\bm{x}}}_{t}^{ll}$, $\hat{{\bm{x}}}_{t}^{lh}$, $\hat{{\bm{x}}}_{t}^{hl}$, and $\hat{{\bm{x}}}_{t}^{hh}$. For simplicity, we perform the same operation for three high frequency components. In order to resist the gradual loss of low-frequency information during the sampling process, we reassign a global adjustment factor $w^{l}_{t}$ greater than 1 to the low-frequency component to enhance the proportion of low-frequency components. In order to cope with the reduction of high-frequency energy at the end of sampling, we assign a coefficient $w^{h}_{t}$ greater than 1 to the high-frequency component.
Figure 5. The overall framework of our method. At each sampling timestep $t$, the predicted noisy image $\hat{{\bm{x}}}_{t}$ is decomposed into four frequency subbands using DWT. Each subband is then assigned a weight coefficient, which is determined based on both the timestep and the type of subband. This weighting is performed via a dot product operation. After the adjustment of each subband, all the modified subbands are combined and restored to the noisy image $\tilde{{\bm{x}}}_{t}$ through iDWT, completing the process of frequency regulation.
(14) $w_{t}^{h}=w_{h}\cdot\mathbbm{1}\{t\leqslant t_{\text{mid}}\}$
Then, we perform inverse wavelet transformations on $w^{l}_{t}\hat{{\bm{x}}}_{t}^{l}$ and $w^{h}_{t}\hat{{\bm{x}}}_{t}^{h}$ to get a new prediction noisy image:
(15) $\tilde{{\bm{x}}}_{t}=\textrm{iDWT}(w_{t}^{f}\hat{{\bm{x}}}_{t}^{f}|f=ll,lh,hl,hh)$
Fig. 6 clearly shows that with the increasing of the weight coefficient $w^{l}_{t}$ of the low-frequency component within a certain range, the basic structure and color distribution of the generated noisy image are gradually clear and prominent, which significantly enhances the visual quality of the image. On the other hand, the change of the high-frequency component weight $w^{h}_{t}$ has a limited impact on the generated image, but the appropriate variation brings about the optimization of texture details to prevent the image from being too smooth due to the loss of high-frequency energy.
(a) $w_{t}^{l}=$1.0
(b) $w_{t}^{l}=$1.002
(c) $w_{t}^{l}=$1.004
(d) $w_{t}^{l}=$1.006
(e) $w_{t}^{h}=$1.0
(f) $w_{t}^{h}=$1.002
(g) $w_{t}^{h}=$1.004
(h) $w_{t}^{h}=$1.006
Figure 6. Effect of frequency component adjustment weight $w_{l}$ and $w_{h}$ on generation quality.
4.4. Weighting Scheme
We find that an excessively large weight for the low-frequency or high-frequency components would cause image distortion, leading to a significant reduction in the quality of the generated noisy images. We speculate that this is because the relatively large weight of frequency components conflicts with the denoising rules of DPM. Therefore, we need to explore more advanced weight strategies to conform to the denoising rules of DPM.
Low frequency weight. DPM focuses on finely restoring the details of the image in the end stage of sampling, as shown in Fig.2. If we still assign a relatively high weight to the low-frequency components of the image at this stage, it may impede the refinement process of the image. This conflict stems from the relative positions of the low-frequency and high-frequency components, and often the weight of the low-frequency components is amplified. To avoid this kind of conflict, we have studied several dynamic weighting strategies that make it decrease over time.
The first is a simple turn-off strategy where the low frequency component is no longer incremented after a certain time step:
(16) $w_{t}^{l}=w_{l}\cdot\mathbbm{1}\{t\geqslant t_{\text{mid}}\},$
where $t_{\text{mid}}$ is a pre-defined time step, which is used to determine when to stop enhancing the low-frequency components. Another one is to utilize the noise variance in the reverse process:
(17) $w_{t}^{l}=1+w_{l}\cdot\sigma_{t}.$
where $\sigma_{t}$ can be either fixed or learned, which is determined by the baseline model. The strategy is based on the variance reflecting the denoising dynamics. Specifically, a larger variance means more reduction to the low-frequency energy of ${\bm{x}}_{t}$ in the forward process, so in the reverse process $\hat{{\bm{x}}}_{t}$ focuses more on restoring low-frequency energy at the current time step. As the variance decreases, the low-frequency energy stabilizes, and the model shifts to high-frequency restoration.
High frequency weight. DPM focus on restoring the low-frequency parts of an image in the early stage of sampling as shown in Fig. 2. In particular, the high-frequency components of the image at this stage are mostly Gaussian noise rather than accurate detailed information because DPM does not restore detailed information at this stage. Therefore, enhancing the high-frequency components at this stage amplifies the Gaussian noise, which obviously has negative effects and hinders the model from restoring the low-frequency energy.
To solve the issue, we propose a closing strategy similar to $w_{t}^{l}$:
(18) $w_{t}^{h}=w_{h}\cdot\mathbbm{1}\{t\leqslant t_{\text{mid}}\},$
where $t_{\text{mid}}$ is consistent with that in Eq. (16). Therefore, $t_{\text{mid}}$ represents the critical time step at which the enhancement of low-frequency components stops and the enhancement of high-frequency components begins. Theoretically, we can also design a weight strategy that increases with the time step based on the noise variance:
(19) $w_{t}^{h}=1+w_{h}\cdot(1-\sigma_{t}),$
corresponding to Eq. (17) of $w_{t}^{l}$. However, we find that the performance of Eq. (19) consistently underperforms Eq. (18). It may stem from the fact high-frequency components in the early stages of sampling predominantly represent noise, which should not be amplified. Our improved algorithm is detailed in Algorithm  1, named W++.
Algorithm 1 W++ Sampling
1:Input: Frequency regulation scaling scale $w^{f}_{t-1}$.
2:${\bm{x}}_{T}\sim\mathcal{N}(0,{\bm{I}})$
3:for $t=T-1,\dots,1,0$ do
4:     ${\bm{z}}\sim\mathcal{N}(0,{\bm{I}})$ if $t>1$, else ${\bm{z}}=0$
5:     ${\bm{x}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)}\right)+\sigma_{t}{\bm{z}}$
6:     ${\bm{x}}_{t-1}^{ll},{\bm{x}}_{t-1}^{lh},{\bm{x}}_{t-1}^{hl},{\bm{x}}_{t-1}^{hh}={\rm\textrm{DWT}}({\bm{x}}_{t-1})$
7:     ${\bm{x}}_{t-1}={\rm\textrm{iDWT}}(w^{f}_{t-1}{\bm{x}}_{t-1}^{f}|f=ll,lh,hl,ll)$
8:end for
9:return ${\bm{x}}_{0}$
5. Experiments
This section presents the experimental setup, including datasets, evaluation metrics, and the implementation details. We provide an analysis of the results, comparing the performance of our method with existing approaches. The experiments demonstrate the effectiveness of our method in improving the generative quality of diffusion models and mitigating exposure bias.
To demonstrate the effectiveness, versatility, and synergy of our methods, we choose a variety of classical diffusion model frameworks, samplers, time steps and datasets for extensive experiments, with various standard metrics, including Fréchet Inception Distance (FID) (Heusel et al., 2017), Inception Score (IS) (Salimans et al., 2016), recall, and precision (Heusel et al., 2017). For all metric reports, 50k generated samples were used, with the complete training set as the reference batch. We emphasize that our research perspective and improved methods are to fill the gap in the study of exposure bias. Therefore, we choose some improved models for exposure bias as the comparison models, such as ADM-ES (Ning et al., 2024) (ICLR 2024), ADM-EP (Li and van der Schaar, 2024) (ICLR 2024), TS-DPM (Li et al., 2024) (ICLR 2024), MDSS (Ren et al., 2024) (AAAI 2024), and La-DDPM (Zhang et al., 2023) (CVPR 2024). Meanwhile, our improved model adds W++ as a suffix, based on a baseline model shown by the prefix.
In this section, we provide the detailed parameter settings for all the experiments in the main text.
Table 2. Settings on CIFAR-10 and Image-Net using ADM.
$w_{l}/w_{h}$ CIFAR-10 Image-Net
$T^{\prime}$ 20 30 50 20 30 50
ADM-W++ $w_{l}$ 1.013 1.008 1.0036 0.050 0.040 0.028
ADM-W++ $w_{h}$ 1.064 1.034 1.015 0.997 0.998 1.001
Table 3. Settings on CIFAR-10 using A-DDPM, EA-DDPM.
$w_{l}/w_{h}$ CIFAR10 (LS) CIFAR10 (CS)
$T^{\prime}$ 10 25 50 10 25 50
A-DDPM $w_{l}$ 0.132 0.049 0.03 0.072 0.032 0.008
A-DDPM $w_{h}$ 1.11 1.038 1.019 1.202 1.046 1.018
NPR-DDPM $w_{l}$ 0.132 0.048 0.024 0.066 0.03 0.007
NPR-DDPM $w_{h}$ 1.105 1.034 1.015 1.192 1.045 1.017
SN-DDPM $w_{l}$ 0.109 0.043 0.025 0.052 0.027 0.009
SN-DDPM $w_{h}$ 1.013 1.005 1.005 1.101 1.019 1.01
Table 4. Settings on CIFAR-10 using DDIM and AMED.
$w_{l}/w_{h}$ DDIM AMED
$T^{\prime}$ 10 25 50 5 7 9
Baseline $w_{l}$ 0.0 0.0 0.0 0.0014 0.0012 0.0027
Baseline $w_{h}$ 1.21 1.037 1.011 0.9955 0.9932 0.9985
Table 5. Settings on CIFAR-10 using EDM and PFGM++.
$w_{l}/w_{h}$ EDM PFGM++
$T^{\prime}$ 13 21 35 13 21 35
Baseline $w_{l}$ 0.036 0.016 0.007 0.037 0.016 0.006
Baseline $w_{h}$ 1.087 1.054 1.022 1.095 1.057 1.025
5.1. Results on ADM, DDPM, and IDDPM
To evaluate the versatility of W++, we take ADM (Dhariwal and Nichol, 2021) as the baseline model, which made DPM surpass GAN by introducing classifier guidance. Meanwhile, we take ADM-ES and ADM-EP as the comparison models, which use train-free sampling techniques to greatly mitigate exposure bias. We conduct unconditional generation on the CIFAR-10 (Krizhevsky et al., 2009) dataset and class-conditional generation on the Image-Net $64\times 64$ (Chrabaszcz et al., 2017) dataset with different timesteps.
Table 6 illustrates ADM-W++ achieves markedly superior performance compared to ADM, ADM-ES, and ADM-EP, regardless of datasets or timesteps. Specifically, in the 20-step and 30-step sampling tasks on CIFAR-10, ADM-W++ attains the remarkable FID, which is less than One half of ADM’s. We observe ADM-W++ exhibits almost excellent performance at any timestep and data set, demonstrating the insensitivity of our approach to experimental settings. Specifically, “-” in tables indicates that the original paper of the comparison models did not report results for the task.
To further compare with the solutions for exposure bias, we once again select DDPM (Ho et al., 2020) and IDDPM (Nichol and Dhariwal, 2021) as the baselines, and TS-DPM and MDSS as the comparison models. Since these two improved models do not use ADM as the baseline, we choose to align with their experimental settings. Table 7 clearly shows that regardless of the timestep or baseline, W++ achieves much higher generation quality than TS-DPM and MDSS.
Table 6. FID on CIFAR-10 and ImageNet using ADM.
CIFAR-10 ImageNet
$T^{\prime}$ 20 30 50 20 30 50
ADM 10.36 6.57 4.43 10.96 6.34 3.75
ADM-EP 7.31 4.95 3.46 - - -
ADM-ES 5.15 3.37 2.49 7.52 4.63 3.07
ADM-W++ 4.56 2.65 2.25 7.18 4.30 2.83
Table 7. FID on CIFAR-10 using DDPM and IDDPM.
Method 10 20 Method 30 100
DDPM 42.04 24.60 IDDPM 7.81 3.72
TS-DPM 33.36 22.21 MDSS 5.25 3.49
DDPM-W++ 13.53 7.48 IDDPM-W++ 3.84 2.77
5.2. Results on A-DDPM and EA-DDPM
To evaluate the versatility of W++, we choose the improved diffusion models A-DDPM (Analytic-DPM) (Bao et al., 2022b) and EA-DDPM (Extended-Analytic-DPM) (Bao et al., 2022a). The former deduced the analytic form of the optimal reverse variance and the latter used neural networks to estimate the optimal covariance of the conditional Gaussian distribution in the inverse process. Specifically, we tested two advanced models in EA-DDPM: NPR-DDPM and SN-DDPM, where the “NPR” and “SN” symbols represent two different methods for estimating optimal variance under different conditions in EA-DDPM. Meanwhile, we conducted various training strategies, employing the linear schedule of $\beta_{n}$ (LS) (Ho et al., 2020), cosine schedule of $\beta_{n}$ (CS) (Nichol and Dhariwal, 2021) respectively. Finally, we chose LA-DDPM (Zhang et al., 2023) as comparison model because this work made improvements on EA-DDPM.
Table 8 demonstrates W++ gets more advanced FID than all baseline models and comparison models regardless of different estimation methods, training strategies, and sampling timesteps, which not only shows W++ can significantly improve generation quality, but also that W++ is more advanced than other improvements. Similar to the pattern observed with ADM, W++ shows the same degree of performance on any timestep and dataset. These results further confirm the effectiveness and versatility of our methods.
Table 8. FID on CIFAR-10 using A-DDPM and EA-DDPM.
CIFAR10 (LS) CIFAR10 (CS)
$T^{\prime}$ 10 25 50 10 25 50
DDPM, $\tilde{\beta}_{n}$ 44.45 21.83 15.21 34.76 16.18 11.11
DDPM, $\beta_{n}$ 233.41 125.05 66.28 205.31 84.71 37.35
A-DDPM 34.26 11.60 7.25 22.94 8.50 5.50
A-DDPM-W++ 12.38 6.63 4.52 11.61 4.40 3.62
NPR-DDPM 32.35 10.55 6.18 19.94 7.99 5.31
LA-NPR-DDPM 25.59 8.84 5.28 - - -
NPR-DDPM-W++ 10.86 5.76 4.11 10.18 4.07 3.44
SN-DDPM 24.06 6.91 4.63 16.33 6.05 4.17
LA-SN-DDPM 19.75 5.92 4.31 - - -
SN-DDPM-W++ 11.73 4.73 3.78 12.53 4.51 3.47
5.3. Results on DDIM, EDM, FPGM++, and AMED
To further verify the effectiveness and versatility of W++, we select several deterministic sampling models DDIM (Song et al., 2021a), EDM (Karras et al., 2022), FPGM++ (Xu et al., 2023), and AMED (Zhou et al., 2024). We choose Analytic-DPM (Bao et al., 2022b) as the backbone model for DDIM and the Euler method as the solver for EDM and PFGM++. Unlike noise-prediction networks, EDM’s neural network restores the original data, PFGM++ is an improved model of the Poisson Flow Generative Model (Xu et al., 2022), and AMED significantly mitigates truncation errors using the median theorem, all of which are crucial for testing the generality of W++. In particular, for DDIM, we use the usual sampling time step as $T^{\prime}$ as in previous experiments, but for EDM, PFGM ++ and AMED, we choose Neural Function Evaluations (NFE) (Vahdat et al., 2021) as $T^{\prime}$.
Table 9 clearly shows that W++ can significantly improve the generation quality of DDIM; For AMED using the mean value theorem and distillation techniques, W++ can still reduce FID by mitigating exposure bias. Although the EDM illustrates the design space of the diffusion models and PFGM++ is a Poisson Flow Generation Model, the results in Table 10 also maintain the same trend, with W++ showing better generation performance than the baseline model and the comparison model. The above experiments show for different deterministic sampling methods, different time steps, and different neural network types, W++ can always significantly improve the generation quality of baseline models, which further proves the effectiveness and versatility of W++.
In all the above experiments, $t_{\rm mid}$ is selected as 0.2.
Table 9. FID on CIFAR-10 using DDIM and AMED.
DDIM AMED
$T^{\prime}$ 10 25 50 5 7 9
Baseline 26.43 9.96 6.02 7.59 4.36 3.67
Baseline-ES 22.64 7.24 4.58 7.53 4.36 3.66
Baseline-W++ 19.50 6.50 4.46 7.01 4.22 3.54
Table 10. FID on CIFAR-10 using EDM and PFGM++.
EDM PFGM++
$T^{\prime}$ 13 21 35 13 21 35
Baseline 10.66 5.91 3.74 12.92 6.53 3.88
Baseline-ES 6.59 3.74 2.59 8.79 4.54 2.91
Baseline-W++ 4.68 2.84 2.13 6.62 3.67 2.53
Table 11. Ablation study of the two regulation schemes.
[…]
ADM 10.55 8.96 0.65 0.53
ADM-low 7.49 9.24 0.66 0.57
ADM-high 7.58 9.67 0.66 0.57
ADM-W++ 4.62 9.74 0.67 0.60
5.4. Qualitative comparison
To visually show the impact of W++ on the generation quality of DPM, we set the same random number seed and sampling steps for ADM and ADM-W++ during sampling to ensure that the sampling trajectories of two models are as similar as possible. Fig. 7 clearly shows that due to exposure bias, the generated samples by ADM exhibit unnatural phenomena such as over-smoothing, over-exposure, or being too dark. However, ADM-W++ significantly alleviates these problems, resulting in more natural-looking generated images. Specifically, ADM-W++ combats the reduction of frequency-domain energy caused by exposure bias by dynamically compensating the low-frequency and high-frequency energy of the image, which further proves the superiority of W++.
Figure 7. Qualitative comparison between ADM (first row) and ADM-W++ (second row) using 20 steps.
5.5. Ablation study
Synergistic Effect. To investigate the synergistic impact of our proposed method, W++, we performed an ablation study using ADM as the baseline. We separately adjusted the low and high-frequency components, denoted as ADM-low and ADM-high, respectively, and then adjusted both components simultaneously, denoted as ADM-W++. For a comprehensive evaluation of fidelity, diversity, and accuracy, we selected Inception Score (IS), recall, and precision as additional metrics. As shown in Table 11, adjusting any single frequency component—either low or high—substantially enhances the generation quality. However, the joint adjustment of both low and high-frequency components consistently outperforms individual adjustments, highlighting the synergistic effect of combining both frequency bands. This result underscores the importance of the simultaneous optimization of both frequency components, as it leads to more balanced and higher-quality generated samples across multiple metrics.
Hyperparameter Insensitivity. To demonstrate the insensitivity of W++ to hyperparameters, specifically the weight coefficients $w_{l}$ for low-frequency components and $w_{h}$ for high-frequency components, we conducted an experiment using A-DDPM and EA-DDPM as baselines. Initially, we adjusted only the low-frequency components of these baselines to find the optimal parameter $w_{l}^{*}$. Then, using the optimal $w_{l}^{*}$, we fine-tuned the high-frequency components to determine the optimal $w_{h}^{*}$. Fig. 8 clearly shows that W++ achieves consistent performance improvements over a wide range of $w_{l}$ and $w_{h}$, suggesting that the method is relatively insensitive to these hyperparameters. This robustness to hyperparameter selection further emphasizes the practical applicability of W++ in real-world scenarios, where manual tuning of such parameters can often be cumbersome and time-consuming.
(a) Search experiments of $w_{l}$
(b) Search experiments of $w_{h}$
Figure 8. Hyperparameters search experiments on CIFAR-10 using A-DDPM and EA-DDPM with 25 steps.
6. Conclusion
In this work, we have addressed the critical challenge of exposure bias in diffusion models, a phenomenon that significantly impairs their generative capabilities. For the first time, we analyze exposure bias through the lens of the frequency domain, revealing that low-frequency energy in the prediction distribution is consistently reduced during the sampling process. Our analytical modeling and extensive experiments confirm this reduction and provide a rigorous mathematical formulation of exposure bias. To mitigate this issue, we introduce a frequency-domain regulation mechanism utilizing wavelet transforms. Our approach effectively preserves low-frequency energy during the early and middle stages of sampling and enhances the prominence of high-frequency details at the final stages. Notably, our method is training-free and plug-and-play, offering significant improvements in the generative quality of various diffusion models. It provides a robust solution to exposure bias, demonstrating the potential for enhanced generation quality across different model architectures.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under the Grant No. 62176108.
Appendix
In this section, we give a detailed proof of the equation 13. We know that the perturbed sample in the forward process is
(20) ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},$
A reconstruction model which predicts ${\bm{x}}_{0}$ given ${\bm{x}}_{t}$:
(21) ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)=\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$
The predicted sample in the reverse process is
(22) $\hat{{\bm{x}}}_{t-1}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{\bm{x}}^{t}_{{\bm{\theta}}}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t}+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1}$
We can substitute Eq. (20) and Eq. (21) into Eq. (22) to obtain:
(23) $\hat{{\bm{x}}}_{t-1}$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}(\gamma_{t}{\bm{x}}_{0}+\phi_{t}{\bm{\epsilon}}_{t})+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}(\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}})$
$+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1}$
$=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}}{1-\bar{\alpha}_{t}}{\bm{x}}_{0}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\phi_{t}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{t}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}{\bm{x}}_{0}$
$+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{1-\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1}$
$=\left(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}}{1-\bar{\alpha}_{t}}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}\right){\bm{x}}_{0}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\phi_{t}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{t}$
$+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{1-\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1}$
We first deal with the first part of Eq. (23):
(24) $\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}}{1-\bar{\alpha}_{t}}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}+\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}$
$=\frac{\sqrt{\bar{\alpha}_{t-1}}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t}}$
$=\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}$
Now, we need to construct an auxiliary term. Therefore, we discard $\gamma_{t}$ in Eq. (24):
(25) $\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}1-\alpha_{t}\bar{\alpha}_{t-1}\big{)}}{1-\bar{\alpha}_{t}}=\sqrt{\bar{\alpha}_{t-1}}$
Since $1-\alpha_{t}>0,\gamma_{t}\leqslant 1$, according to Eq. (24) and Eq. (25), we can naturally obtain:
(26) $\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}\leqslant\sqrt{\bar{\alpha}_{t-1}}$
We can definitely define a new coefficient $\gamma_{t-1}\leqslant 1$ such that:
(27) $\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}$
For the second part of Eq. (23), we can always split it into:
(28) $Var(\hat{{\bm{x}}}_{t-1})$ $=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\sqrt{1-\bar{\alpha}_{t-1}})^{2}+\tilde{\beta}_{t}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\sqrt{1-\bar{\alpha}_{t-1}})^{2}$
$+\frac{(1-\bar{\alpha}_{t-1})(1-\alpha_{t})}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{\alpha_{t}(1-\bar{\alpha}_{t-1})^{2}}{1-\bar{\alpha}_{t}}+\frac{(1-\bar{\alpha}_{t-1})(1-\alpha_{t})}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{\alpha_{t}(1-\bar{\alpha}_{t-1})^{2}+(1-\bar{\alpha}_{t-1})(1-\alpha_{t})}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{(1-\bar{\alpha}_{t-1})[\alpha_{t}(1-\bar{\alpha}_{t-1})+(1-\alpha_{t})]}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{(1-\bar{\alpha}_{t-1})[\alpha_{t}-\bar{\alpha}_{t}+1-\alpha_{t}]}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{(1-\bar{\alpha}_{t-1})[1-\bar{\alpha}_{t}]}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+1-\bar{\alpha}_{t-1}$
In order to match the mean term, we rewrite Eq. (28) as:
(29) $Var(\hat{{\bm{x}}}_{t-1})=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(1-\gamma_{t}^{2})(1-\bar{\alpha}_{t-1})+\gamma_{t}^{2}(1-\bar{\alpha}_{t-1})$
Finally, we can obtain:
(30) $\bm{\mu}\big{(}q_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t-1}|{\bm{x}}_{t})\big{)}=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}{\bm{x}}_{0},\quad Var\big{(}q_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t-1}|{\bm{x}}_{t})\big{)}=$
$\gamma_{t}^{2}(1-\bar{\alpha}_{t-1})+(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(1-\gamma_{t}^{2})(1-\bar{\alpha}_{t-1})$
According to Eq. (20), Eq. (30) is equivalent to:
(31) $\bm{\mu}\big{(}q_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t-1}|{\bm{x}}_{t})\big{)}=\gamma_{t-1}{\bm{x}}_{t-1},\quad Var\big{(}q_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t-1}|{\bm{x}}_{t})\big{)}=$
$(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(1-\gamma_{t}^{2})(1-\bar{\alpha}_{t-1})$
In summary, we can obtain:
(32) $\hat{{\bm{x}}}_{t-1}$ $=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}{\bm{x}}_{0}+$
$\sqrt{\gamma_{t}^{2}(1-\bar{\alpha}_{t-1})+(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(1-\gamma_{t}^{2})(1-\bar{\alpha}_{t-1})}{\bm{\epsilon}}_{t-1}$
$=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t-1}+(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}}{\bm{\epsilon}}_{t-1}$