### v2

# Frequency Regulation for Exposure Bias Mitigation in Diffusion Models

Abstract.
Diffusion models exhibit impressive generative capabilities but are significantly impacted by exposure bias. In this paper, we make a key observation: the energy of predicted noisy samples in the reverse process continuously declines compared to perturbed samples in the forward process. Building on this, we identify two important findings: 1) The reduction in energy follows distinct patterns in the low-frequency and high-frequency subbands; 2) The subband energy of reverse-process reconstructed samples is consistently lower than that of forward-process ones, and both are lower than the original data samples. Based on the first finding, we introduce a dynamic frequency regulation mechanism utilizing wavelet transforms, which separately adjusts the low- and high-frequency subbands. Leveraging the second insight, we derive the rigorous mathematical form of exposure bias. It is worth noting that, our method is training-free and plug-and-play, significantly improving the generative quality of various diffusion models and frameworks with negligible computational cost. The source code is available at https://github.com/kunzhan/wpp.
Diffusion Model, Exposure Bias, Wavelet Transform
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755608submissionid: 4742ccs: Computing methodologies Computer vision
1. Introduction
In recent years, Diffusion Probabilistic Models (DPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) have made remarkable progress in image generation (Dhariwal and Nichol, 2021; Rombach et al., 2022). ADM (Dhariwal and Nichol, 2021) made the generation quality of DPMs surpass Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) by introducing classifier guidance. EDM (Karras et al., 2022) notably promotes development by clarifying the design space of DPMs. Additionally, IDDPM (Nichol and Dhariwal, 2021), DDIM (Song et al., 2021a), Analytic-DPM (Bao et al., 2022b), EA-DPM (Bao et al., 2022a), PFGM++ (Xu et al., 2023), and AMED (Zhou et al., 2024) also promote the improvement of DPMs from different aspects. However, these models still suffer from exposure bias (Ning et al., 2023), the mismatch between the forward and reverse process in DPMs. Due to the prediction error (Kim et al., 2023) of the network and the discretization error (Zhang and Chen, 2023) of the numerical solver, the reverse trajectory of DPMs tends to deviate from the ideal path. Meanwhile, the bias will gradually accumulate during sampling, ultimately affecting the generation quality.
(a) The low-frequency energy.
(b) The high-frequency energy.
Figure 1. The subband energy of the noisy sample in training and sampling, with denoising from right to left.
A variety of methods have been proposed to mitigate exposure bias. These approaches include diverse training strategies (Ning et al., 2023; Li and van der Schaar, 2024; Ren et al., 2024) and sampling techniques (Li et al., 2024; Zhang et al., 2025; YAO et al., 2025). However, all these methods are confined to operating in the spatial domain, neglecting the analysis and mitigation of exposure bias in the wavelet domain. We observed during the reverse process, the energy of the predicted noisy sample is always lower than that of the forward perturbed sample. Interestingly, this reduction pattern exhibits distinctly different modes in different frequency subbands. By applying Discrete Wavelet Transform (DWT), we analyze the evolution of exposure bias within the wavelet domain and uncover the first key findings: low-frequency subband energy reduction persists throughout sampling, with high-frequency energy primarily diminishing in the later reverse stage, as shown in Fig. 1. Applying the same method, we analyze the reconstructed sample, which directly predicts the original sample given the noisy sample. Thus we have the second key observation: the subband energy of reverse-process reconstructed samples is consistently lower than that of forward-process ones, and both are lower than the original data samples.
Based on the first finding, we propose a simple yet effective frequency-domain information adjustment mechanism. The predicted noisy sample is decomposed into high- and low-frequency subbands via DWT. Then, the low-frequency subband is amplified with a dedicated weight throughout sampling, while the high-frequency subbands are amplified in the late sampling using the high-frequency weight. Finally, the adjusted subbands are combined and reconstructed into the predicted noisy sample in the spatial space via inverse Discrete Wavelet Transform.
Building on the second finding, we derive the mathematical analytical form of exposure bias. Previous work has severely lacked exploration into the analytical form of exposure bias. In particular, their modeling of exposure bias relies heavily on strong assumptions (Ning et al., 2024; Li et al., 2024). Leveraging the discovery of the energy law of reconstructed samples, we conduct more accurate modeling and derivation, providing a clearer understanding of exposure bias.
Additionally, we observe amplifying the low-frequency subband towards the end of the sampling process may conflict with the denoising principles of DPMs, potentially introducing undesirable effects. Specifically, DPMs aim to finely restore image details in the later stages of sampling. As shown in Fig. 2, high-frequency information begins to change significantly only at the final stages. If a relatively large weight is assigned to the low-frequency subband during this phase, it hinders the refinement process. To address this issue, we explore several dynamic weighting strategies, where the amplifying effect of the low-frequency subband decreases as sampling progresses, while that of high-frequency components gradually increases. In summary, our contributions are:
[…]
  • We are, to the best of our knowledge, the first to analyze and solve the exposure bias of DPMs from the perspective of the wavelet domain. Furthermore, based on findings in the wavelet domain, we derive for the first time the rigorous mathematical form of exposure bias.
  • We propose a dynamic frequency regulation mechanism based on wavelet transform to mitigate exposure bias. The mechanism is train-free and plug-and-play, improving the generation quality of various DPMs.
  • Our method can achieve high-quality performance metrics with negligible computational cost, outperforming all current improved models for exposure bias.
2. Related Work
This section first reviews the development of DPMs, then presents some work on improving DPMs using wavelet transforms, and finally introduces existing methods for mitigating exposure bias.
The theoretical foundation of DPMs was established by DPM (Sohl-Dickstein et al., 2015), with substantial advancements later achieved by DDPM (Ho et al., 2020). By introducing classifier guidance, ADM (Dhariwal and Nichol, 2021) made the generation performance of DPMs surpass GAN (Goodfellow et al., 2014) for the first time. SDE (Song et al., 2021b) proposes a unified DPM framework by means of stochastic differential equations, while DDIM (Song et al., 2021a) accelerates sampling by skipping steps. Analytic-DPM (Bao et al., 2022b) theoretically derived the optimal variance in the reverse denoising process, while EDM (Karras et al., 2022) clarified the design space of DPMs, which further enhances the generation quality of DPMs. In addition, ODE-based DPMs (Lu et al., 2022; Zhou et al., 2024; Zhao et al., 2024; zhangfast; Dockhorn et al., 2022), DPMs incorporating distillation techniques (Salimans and Ho, 2022; Liu et al., 2023; Meng et al., 2023; Luhman and Luhman, 2021; Zheng et al., 2023), and consistency models (Song et al., 2023; Song and Dhariwal, 2024; Lu and Song, 2025) are widely developed, which significantly speeds up sampling and improves the generation quality.
Wavelet decomposition (Graps, 1995; Mallat, 1989) has been widely applied in the field of GANs (Gal et al., 2021; Yang et al., 2022; Wang et al., 2022; Zhang et al., 2022), improving the quality of various generation tasks. In recent years, some works have started to combine wavelet transform with DPMs. WACM  (Li et al., 2022) utilizes the wavelet spectrum to assist image colorization. WSGM (Guth et al., 2022) factorizes the data distribution into the product of conditional probabilities of wavelet coefficients at various scales to accelerate the generation process. WaveDiff (Phung et al., 2023) processes images in the wavelet domain to further enhance the efficiency and quality of generation.
The exposure bias in DPMs was first systematically analyzed by ADM-IP (Ning et al., 2023), which simulated the bias by re-perturbing the training sample. Subsequently, EP-DDPM (Li and van der Schaar, 2024) regard the cumulative error as the regularization term to retrain the model for alleviating exposure bias, while MDSS (Ren et al., 2024) employed the multi-step denoising scheduled sampling strategy to mitigate this issue. It should be noted these three methods require retraining the model. Conversely, TS-DPM (Li et al., 2024) put forward the time-shift sampler and ADM-ES applied the noise scaling technique, which both can mitigate exposure bias without retraining models. Additionally, AE-DPM (Zhang et al., 2025) alleviates the bias through prompt learning, and MCDO (YAO et al., 2025) mitigates it by means of manifold constraint. S++ (Yu and Zhan, 2025) proposes a clever score difference correction mechanism to reduce exposure bias.
Unfortunately, all previous studies on exposure bias have been confined to the spatial domain, also called the pixel domain. In contrast, we analyze this issue from the perspective of the wavelet domain and address it using frequency-domain approaches.
3. Preliminaries
In this section, we introduce the necessary fundamental knowledge required to understand our problem and approach. We first introduce the Diffusion Probabilistic Model (DPM) (Ho et al., 2020), followed by the exposure bias (Ning et al., 2023) and the wavelet transform.
3.1. Diffusion model
Diffusion Probabilistic Model (DPM) usually consists of a forward noising process and a reverse denoising process, both of which are defined as Markov processes. For a original data distribution $q({\bm{x}}_{0})$ and a noise schedule $\beta_{t}$, the forward process is
(1) $q\left({\bm{x}}_{1:T}|{\bm{x}}_{0}\right)=\prod_{t=1}^{T}q\left({\bm{x}}_{t}|{\bm{x}}_{t-1}\right),$
where $q\left({\bm{x}}_{t}|{\bm{x}}_{t-1}\right)=\mathcal{N}({\bm{x}}_{t};\sqrt{1-\beta_{t}}{\bm{x}}_{t-1},\beta_{t}{\bm{I}})$. Specifically, by leveraging the properties of the Gaussian distribution, the perturbed distribution can be rewritten as the conditional distribution $q\left({\bm{x}}_{t}|{\bm{x}}_{0}\right)$, which is expressed by the formula
(2) ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{t},$
where ${\alpha}_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and ${\bm{\epsilon}}_{t}\sim\mathcal{N}(\bm{0},{\bm{I}})$. Based on Bayes’ theorem, the posterior conditional probability is obtained, which is
(3) $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})=\mathcal{N}(\tilde{\mu}_{t}({\bm{x}}_{t},{\bm{x}}_{0}),\tilde{\beta}_{t}{\bm{I}}),$
where $\tilde{\mu}_{t}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}{\bm{x}}_{0}+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}{\bm{x}}_{t}$ and $\tilde{\beta}_{t}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}$. We apply a neural network $p_{{{\bm{\theta}}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}({\bm{x}}_{t-1};\mu_{{\bm{\theta}}}({\bm{x}}_{t},t),\sigma_{t}{\bm{I}})$ to approximate $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$, and naturally we want to minimize $D_{\rm KL}(q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||p_{{\bm{\theta}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})))$. Through a series of reparameterizations and derivations, we obtain
(4) $\mu_{{\bm{\theta}}}({\bm{x}}_{t},t)$ $=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}{\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}{\bm{x}}_{t}$
$=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)\right),$
where ${\bm{\epsilon}}_{\bm{\theta}}(\cdot)$ is the noise prediction network and ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ represent the reconstruction model which predicts ${\bm{x}}_{0}$ given ${\bm{x}}_{t}$:
(5) ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)=\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$
So the final loss function is given by
(6) $\mathcal{L}_{\rm simple}=\mathbb{E}_{t,{\bm{x}}_{0},{\bm{\epsilon}}_{t}\sim\mathcal{N}(\mathbf{0},{\bm{I}})}[\|{\bm{\epsilon}}_{{\bm{\theta}}}({\bm{x}}_{t},t)-{\bm{\epsilon}}_{t}\|_{2}^{2}].$
Once ${\bm{\epsilon}}_{{\bm{\theta}}}(\cdot)$ has converged, we can perform reverse sampling starting from a standard Gaussian distribution by using $p_{{{\bm{\theta}}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})$.
3.2. Exposure bias
The exposure bias is the mismatch between the forward and reverse processes in DPMs. Specifically, for the pre-trained noise prediction network ${\bm{\epsilon}}_{\bm{\theta}}(\cdot)$, the network input comes from the perturbed noisy sample ${\bm{x}}_{t}$ based on the ground truth during training, while during sampling the network input comes from the predicted noisy sample $\hat{{\bm{x}}}_{t}$. Due to the prediction error of the neural network and the discretization error of the numerical solver, the sampling trajectory of DPM tends to deviate from the ideal path, which leads to the bias between the predicted noisy sample $\hat{{\bm{x}}}_{t}$ and the perturbed noisy sample ${\bm{x}}_{t}$. This exposure bias between noisy samples leads to a bias in the network output ${\bm{\epsilon}}_{\bm{\theta}}\left(\hat{{\bm{x}}}_{t},t\right)$ and ${\bm{\epsilon}}_{\bm{\theta}}\left({\bm{x}}_{t},t\right)$, which exacerbates the exposure bias in the next time step. Therefore, the exposure bias gradually accumulates with the sampling iteration, and ultimately affect the generation quality, as shown in Fig. 3. Thus, due to exposure bias, the actual sampling formula should be
Figure 3. The schematic diagram of exposure bias in DPM.
(7) $\hat{{\bm{x}}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\hat{{\bm{x}}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)\right)+\sigma_{t}{\bm{z}}.$
Specially, both $\hat{{\bm{x}}}_{T}$ and ${\bm{x}}_{T}$ obey the standard Gaussian distribution. However, due to the prediction error between ${\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{T},T)$ and the ideal noise, the sampling operation in Eq. (7) leads to the deviation between $\hat{{\bm{x}}}_{T-1}$ and ${\bm{x}}_{T-1}$, which further exacerbates the bias between ${\bm{\epsilon}}_{\bm{\theta}}(\hat{{\bm{x}}}_{T-1},T-1)$ and ${\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{T-1},T-1)$. Thus, as sampling continues, the bias of the predicted noisy sample and the bias of the network prediction influences each other and gradually accumulate, causing the exposure bias between $\hat{{\bm{x}}}_{t}$ and ${\bm{x}}_{t}$.
3.3. Wavelet transform
We introduce Discrete Wavelet Transform (DWT), which decomposes an image ${\bm{x}}\in\mathbb{R}^{H\times W}$ into four wavelet subbands, namely ${\bm{x}}^{ll},{\bm{x}}^{lh},{\bm{x}}^{hl}$, and ${\bm{x}}^{hh}$. Conversely, these wavelet subbands can be reconstructed into an image by using the inverse Discrete Wavelet Transform (iDWT). The two processes are expressed as:
(8) $\begin{cases}\{{\bm{x}}^{f}|f\in\{ll,lh,hl,hh\}\}=\textrm{DWT}({\bm{x}})\\ {\bm{x}}=\textrm{iDWT}({\bm{x}}^{f}|f\in\{ll,lh,hl,hh\})\end{cases}$
where the size of all four subbands is $\mathbb{R}^{H/2\times W/2}$. ${\bm{x}}^{ll}$ represents the low-frequency component of the image ${\bm{x}}$, reflecting the basic structure of the image, such as the human face shape and the shape of birds. ${\bm{x}}^{lh},{\bm{x}}^{hl}$, and ${\bm{x}}^{hh}$ represent the high-frequency components in different directions of the image, reflecting the detailed information of the image, such as human wrinkles and bird feathers. For simplicity, we choose the classic Haar wavelet in experiments.
4. methods
We first analyze exposure bias in the wavelet domain using Discrete Wavelet Transform (DWT). the key finding reveal that the low-frequency subband energy of predicted samples in the reverse process is consistently lower than that of perturbed samples in the forward process throughout the entire sampling process, while the high-frequency subband energy decreases only at the end of sampling. Then, we similarly use DWT to analyze the energy evolution patterns of reconstructed samples in both forward and reverse processes, based on which a reasonable hypothesis is proposed to derive the analytical form of exposure bias. Finally, a simple yet effective frequency-energy adjustment mechanism is proposed to mitigate exposure bias, which works by adjusting the frequency-energy distribution at each time step of the sampling process. Additionally, considering the rule that the reverse process of DPMs first reconstructs low-frequency information and then focuses on recovering high-frequency, a novel dynamic weighting strategy is proposed to further precisely scale the energy of frequency subbands.
4.1. Energy reduction in frequency space
(a) The $ll$ subband of ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$
(b) The $lh$ subband of ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$
Figure 4. The subband energy of ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$ in training and sampling, with denoising from right to left.
Here, we analyze exposure bias in the wavelet domain. Firstly, we need to simulate exposure bias accurately. During sampling, it is easy to obtain the predicted noisy $\hat{{\bm{x}}}_{t}$, but acquiring the ideal perturbed noisy ${\bm{x}}_{t}$ corresponding to $\hat{{\bm{x}}}_{t}$ is difficult. To solve this problem, we rely on the original data and the deterministic solver to simulate the exposure bias. Specifically, for the given original data ${\bm{x}}_{0}$, we first perform forward perturbation on ${\bm{x}}_{0}$ to obtain a series of perturbed noisy $\{{\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{s+1}\}$. Then, we take ${\bm{x}}_{s+1}$ as the actual sampling starting point, match the corresponding time step $s+1$ and coefficients, and carry out the DDIM sampling (Song et al., 2021a)
(9) $\hat{{\bm{x}}}_{s-1}=\sqrt{\bar{\alpha}_{s-1}}\left(\frac{x_{s}-\sqrt{\bar{\alpha}_{s}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{s},s)}{\sqrt{\bar{\alpha}_{s}}}\right)+\sqrt{1-\bar{\alpha}_{s-1}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{s},s).$
As a result, a series of predicted noisy sample $\{\hat{{\bm{x}}}_{s-1},\ldots,\hat{{\bm{x}}}_{2},\hat{{\bm{x}}}_{1},\hat{{\bm{x}}}_{0}\}$ is obtained. As long as we ensure that ${\bm{x}}_{s}$ still maintain a certain level of low-frequency information of original data ${\bm{x}}_{0}$, we are able to make the reverse predicted noisy sample $\hat{{\bm{x}}}_{t}$ similar to the forward perturbed noisy sample ${\bm{x}}_{t}$.
Secondly, we perform Discrete Wavelet Transform (DWT) on the perturbed and predicted noisy sample at different time steps to get their corresponding frequency components. Then, we calculate respectively the mean squared norm of the subbands of these two batches of noisy sample at different time steps. Fig. 1(a) clearly shows that as the sampling progresses, the energy of the low-frequency subbands of $\hat{{\bm{x}}}_{t}$ gradually becomes lower than that of ${\bm{x}}_{t}$, which indicates there is a phenomenon of reducing of low-frequency energy during sampling. Fig. 1(b) shows only at the end of sampling does the high-frequency subband energy of $\hat{{\bm{x}}}_{t}$ become significantly lower than that of ${{\bm{x}}}_{t}$ (during the early and middle stages of sampling, the two are comparable). Intuitively, the diffusion model starts to reconstruct low-frequency information in the early stage of sampling, so the deviation of the low-frequency subband will accumulate and propagate throughout the entire sampling process. In contrast, high-frequency information is only restored emphatically at the end stage, so the deviation of the high-frequency subband is only manifested at the end stage of sampling.
The above experiments indicate there is a reduction of low-frequency energy throughout the sampling process and a reduction of high-frequency energy at the end of sampling, which is the manifestation of exposure bias in the wavelet domain.
Table 1. Mean and variance of $q({\bm{x}}_{t}|{\bm{x}}_{0})$ and $p_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t}|{\bm{x}}_{t+1})$
Mean Variance
$q({\bm{x}}_{t}|{\bm{x}}_{0})$ $\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}$ $(1-\bar{\alpha}_{t})\bm{I}$
$p_{{\bm{\theta}}}(\hat{{\bm{x}}}_{t}|{\bm{x}}_{t+1})$ $\gamma_{t}\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}$ $\left(1-\bar{\alpha}_{t}+\left(\frac{\sqrt{\bar{\alpha}_{t}}\beta_{t+1}}{1-\bar{\alpha}_{t+1}}\phi_{t+1}\right)^{2}\right)\bm{I}$
4.2. Theoretical perspective
Then, we use the same method to explore the bias pattern of reconstructed samples and derive the analytical form of exposure bias. Previous work (Ning et al., 2024; Li et al., 2024) rely heavily on a strong assumption
(10) ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)={\bm{x}}_{0}+\phi_{t}{\bm{\epsilon}}_{t},$
where ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ represent the reconstruction model predicting ${\bm{x}}_{0}$ given ${\bm{x}}_{t}$, shown in Eq. (5), and $\phi_{t}$ is a scalar coefficient related to time steps. However, we will overthrow this assumption.
To explore the modeling rule of ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$, we used Eq. (5) to obtain ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ and ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)$ corresponding to ${\bm{x}}_{t}$ and $\hat{{\bm{x}}}_{t}$ based on the experiment in §4.1, and performed the same wavelet decomposition and norm calculation Figs. 4(a) and 4(b) clearly show that neither ${\bm{x}}_{t}$ nor $\hat{{\bm{x}}}_{t}$ can fully reconstruct ${\bm{x}}_{0}$ at any time step, and ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)$ reconstructed by the predicted noisy sample is smaller than ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)$ reconstructed by the perturb noisy sample regardless of the frequency component and time step. Based on this key observation, we make a more scientific assumption about ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$.
Assumption 1 Whether in the forward process or the reverse process, ${\bm{x}}^{0}_{\bm{\theta}}({{\bm{x}}}_{t},t)$ is assumed to be represented in terms of ${\bm{x}}_{0}$ as
(11) ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)=\gamma_{t}{\bm{x}}_{0}+\phi_{t}{\bm{\epsilon}}_{t},$
where ${\bm{\epsilon}}_{t}\sim\mathcal{N}(\bm{0},{\bm{I}})$, $0<\gamma_{t}\leqslant 1$, $\phi_{t}<M$, and $M$ is an upper bound. Based on this assumption, we re-parameterize the conditional probability in the sampling process:
(12) $\hat{{\bm{x}}}_{t-1}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}{\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}{\bm{x}}_{t}+\sqrt{\tilde{\beta_{t}}}{\bm{\epsilon}}_{1}.$
Through plugging Eqs. (11) and (2) into Eq. (12), we obtain the analytical form of $\hat{{\bm{x}}}_{t}$:
(13) $\hat{{\bm{x}}}_{t-1}=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t-1}+\left(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t}\right)^{2}}{{\bm{\epsilon}}}_{2},$
where ${\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$. And the proof of  13 is provided in the appendix. Table 1 clearly shows that the predicted distribution always has a larger variance and a smaller mean than the perturbation distribution, indicating that prediction noisy sample reduces energy about the target data ${\bm{x}}_{0}$ during sampling. Naturally, as the diffusion model focuses on recovering low-frequency energy in early stages, predictions mainly lose low-frequency components then; in late stages, when the model reconstructs high-frequency details, predictions primarily lose high-frequency components.
4.3. Frequency regulation
Based on previous key findings, there are different patterns of reduction in the frequency subband energies of the predicted samples. Naturally, we hope to enhance low-frequency subband energy throughout sampling process and high-frequency subband energy at its end. Thus, we propose a simple and effective mechanism for wavelet domain component regulation based on Discrete Wavelet Transform (DWT), as shown in Fig. 5. Specifically, for the predicted noisy sample $\hat{{\bm{x}}}_{t}$ at each step in sampling process, we use the DWT to decompose it into four distinct frequency components: $\hat{{\bm{x}}}_{t}^{ll}$, $\hat{{\bm{x}}}_{t}^{lh}$, $\hat{{\bm{x}}}_{t}^{hl}$, and $\hat{{\bm{x}}}_{t}^{hh}$. For simplicity, we perform the same operation for three high frequency components. In order to resist the gradual loss of low-frequency information during the entire sampling process, we reassign a global adjustment factor $w^{l}_{t}$ greater than 111 to the low-frequency component to enhance the proportion of low-frequency component. In order to cope with the reduction of high-frequency energy at the end of sampling, we also assign a coefficient $w^{h}_{t}$ greater than 111 to the high-frequency component. However, this value is related to the sampling process:
Figure 5. The overall framework of our method. At each sampling timestep $t$, the predicted noisy sample $\hat{{\bm{x}}}_{t}$ is decomposed into four frequency subbands using DWT. Each subband is then assigned a weight coefficient, which is determined based on both the timestep and the type of subband. This weighting is performed via a dot product operation. After the adjustment of each subband, all the modified subbands are combined and restored to the noisy sample $\tilde{{\bm{x}}}_{t}$ through iDWT, completing the process of frequency regulation.
(14) $w_{t}^{h}=w_{h}\cdot\mathbbm{1}\{t\leqslant t_{\text{mid}}\}$
Then, we perform inverse Discrete Wavelet Transform (iDWT) on the adjusted frequency subbands to get the new adjusted sample:
(15) $\tilde{{\bm{x}}}_{t}=\textrm{iDWT}(w_{t}^{f}\hat{{\bm{x}}}_{t}^{f}|f\in\{ll,lh,hl,hh\})$
Fig. 6 clearly shows that with the increasing of the weight coefficient $w^{l}_{t}$ of the low-frequency component within a certain range, the basic structure and color distribution of the generated noisy sample are gradually clear and prominent, which significantly enhances the visual quality of the image. On the other hand, the change of the high-frequency component weight $w^{h}_{t}$ has a limited impact on the generated image, but the appropriate variation brings about the optimization of texture details to prevent the image from being too smooth due to the loss of high-frequency energy.
(a) $w_{t}^{l}=$1.0
(b) $w_{t}^{l}=$1.002
(c) $w_{t}^{l}=$1.004
(d) $w_{t}^{l}=$1.006
(e) $w_{t}^{h}=$1.0
(f) $w_{t}^{h}=$1.002
(g) $w_{t}^{h}=$1.004
(h) $w_{t}^{h}=$1.006
Figure 6. Effect of frequency component adjustment weight $w_{l}$ and $w_{h}$ on generation quality.
4.4. Weighting Scheme
We find that an excessively large weight for the low-frequency or high-frequency components would cause image distortion, leading to a significant reduction in the quality of the generated sample. We speculate that this is because the relatively large weight of frequency components conflicts with the denoising rules of DPM. Therefore, we need to explore more advanced weight strategies to conform to the denoising rules of DPMs.
Low frequency weight. DPMs focus on finely restoring the details of the image in the end stage of sampling process, as shown in Fig 2. Thus, if we still assign a relatively high weight to the low-frequency component of the image at this stage, it may impede the refinement process of the image. This conflict arises from the relative interaction of low- and high-frequency components, and over-amplifying the former will weaken the latter. To avoid this kind of conflict, we have studied several dynamic weighting strategies that make it decrease as sampling proceeds.
The first is a simple turn-off strategy where the low frequency component is no longer enhanced after a critical time step:
(16) $w_{t}^{l}=w_{l}\cdot\mathbbm{1}\{t\geqslant t_{\text{mid}}\},$
where $t_{\text{mid}}$ is a pre-defined time step, which is used to determine when to stop enhancing the low-frequency component. Another one is to utilize the variance in the reverse process:
(17) $w_{t}^{l}=1+w_{l}\cdot\sigma_{t},$
where $\sigma_{t}$ can be either fixed or learned, which is determined by the baseline model. The strategy is based on the variance reflecting the denoising process. Specifically, a larger variance means more reduction to the low-frequency component of ${\bm{x}}_{t}$ in the forward process, thus in the reverse process, DPMs should focus more on restoring the low-frequency component at the current timestep. As the variance decreases, the low-frequency component stabilizes, and DPMs shift to high-frequency components restoration.
High frequency weight. DPMs focus on restoring the low-frequency part in the early stage of sampling as shown in Fig. 2. In particular, high-frequency components of the image at this stage are mostly Gaussian noise rather than accurate detailed information because DPMs do not restore useful high-frequency information at this stage. Therefore, enhancing high-frequency components at this stage will amplify the noise, which obviously has negative effects and hinders DPMs from restoring the low-frequency information. To solve the issue, we propose a turn-off strategy similar to $w_{t}^{l}$:
(18) $w_{t}^{h}=w_{h}\cdot\mathbbm{1}\{t\leqslant t_{\text{mid}}\},$
where $t_{\text{mid}}$ is consistent with that in Eq. (16). Therefore, $t_{\text{mid}}$ represents the critical timestep at which the enhancement of the low-frequency component stops and the enhancement of high-frequency components begins. We can also design a weight strategy that increases with the time step based on the variance:
(19) $w_{t}^{h}=1+w_{h}\cdot(1-\sigma_{t}),$
corresponding to Eq. (17) of the low-frequency component. However, we find that the performance of Eq. (19) consistently underperforms Eq. (18). It may stem from the fact high-frequency components in the early stages of sampling predominantly represent noise, which should not be amplified. Finally, our improved model adds W++ as a suffix, based on a baseline model shown by the prefix, with the algorithm detailed in Alg 1.
Algorithm 1 W++ Sampling.
1:Input: Frequency regulation scaling scale $w^{f}_{t-1}$
[…]
3:for $t=T-1,\dots,1,0$ do
4:  ${\bm{z}}\sim\mathcal{N}(0,{\bm{I}})$ if $t>1$, else ${\bm{z}}=0$
5:  ${\bm{x}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)}\right)+\sigma_{t}{\bm{z}}$
6:  ${\bm{x}}_{t-1}^{ll},{\bm{x}}_{t-1}^{lh},{\bm{x}}_{t-1}^{hl},{\bm{x}}_{t-1}^{hh}={\rm\textrm{DWT}}({\bm{x}}_{t-1})$
7:  ${\bm{x}}_{t-1}={\rm\textrm{iDWT}}\big{(}w^{f}_{t-1}{\bm{x}}_{t-1}^{f}|f\in\{ll,lh,hl,ll\}\big{)}$
8:end for
9:return ${\bm{x}}_{0}$
5. Experiments
In this section, we conduct extensive experimental tests and present detailed experimental setups, test results, experimental analyses, and ablation comparisons. Meanwhile, we introduce a large number of improved models targeting exposure bias for comparison. We emphasize that our method can significantly mitigate exposure bias and thereby improve the generation quality of DPMs.
To demonstrate the effectiveness, versatility, and practicality of our methods, we choose a variety of classical diffusion model frameworks and samplers, such as DDPM (Ho et al., 2020), IDDPM (Nichol and Dhariwal, 2021), ADM (Dhariwal and Nichol, 2021), DDIM (Song et al., 2021a), A-DPM (Bao et al., 2022b), EA-DPM (Bao et al., 2022a), EDM (Karras et al., 2022), PFGM++ (Xu et al., 2023), and AMED (Zhou et al., 2024). Without loss of generality, datasets with different resolutions are selected, such as CIFAR-10 (Krizhevsky et al., 2009), ImageNet $\times 64$ 64, ImageNet $\times 128$ 128 (Chrabaszcz et al., 2017), and LSUN bedroom $\times 256$ 256 (Yu et al., 2015). In general, our experiments are divided into two categories: stochastic generation (Ho et al., 2020) and deterministic generation (Song et al., 2021b). Meanwhile, we select different sampling steps for each type of experiment. To more accurately evaluate the generation quality, we select a variety of evaluation metrics, including Fréchet Inception Distance (FID) (Heusel et al., 2017), Inception Score (IS) (Salimans et al., 2016), recall, and precision (Heusel et al., 2017). For all metric reports, 50k generated samples were used, with the complete training set as the reference batch. To more intuitively demonstrate the effectiveness of method, we also present qualitative displays of the generated images. Specially, we emphasize our research perspective and improved methods are to fill the gap in the study of exposure bias. Thus, we choose some improved models for exposure bias as the comparison models, such as La-DDPM (Zhang et al., 2023) (CVPR 2023), ADM-IP (Li and van der Schaar, 2024) (ICML 2023), MDSS (Ren et al., 2024) (AAAI 2024) ADM-ES (Ning et al., 2024) (ICLR 2024), and TS-DPM (Li et al., 2024) (ICLR 2024). To demonstrate the synergy and robustness of our method, we have conducted a large number of ablation experiments, including comparative analysis between single methods and the integrated method, hyperparameter sensitivity tests, and evaluation of computational resource consumption.
Table 2. FID on CIFAR-10 and ImageNet using ADM.
CIFAR-10 ImageNet
$T^{\prime}$ 20 30 50 20 30 50
ADM 10.36 6.57 4.43 10.96 6.34 3.75
ADM-IP 6.89 4.25 2.92 - - -
ADM-ES 5.15 3.37 2.49 7.52 4.63 3.07
ADM-W++ 4.56 2.65 2.25 7.18 4.30 2.83
Table 3. FID on CIFAR-10 using DDPM and IDDPM.
Method 10 20 Method 30 100
DDPM 42.04 24.60 IDDPM 7.81 3.72
TS-DPM 33.36 22.21 MDSS 5.25 3.49
DDPM-W++ 13.54 7.48 IDDPM-W++ 3.84 2.77
5.1. Results on ADM, DDPM, and IDDPM
To evaluate the versatility of W++, we take ADM (Dhariwal and Nichol, 2021) as the baseline model, which made DPMs surpass GANs by introducing classifier guidance. Meanwhile, we take ADM-IP (Ning et al., 2023) and ADM-ES (Ning et al., 2024) as the comparison models, which mitigated exposure bias from the two perspectives of training and sampling, respectively. We conduct unconditional generation on the CIFAR-10 $\times 32$ 32 (Krizhevsky et al., 2009) dataset and class-conditional generation on the Image-Net $\times 64$ 64 (Chrabaszcz et al., 2017) dataset with different sampling timesteps.
Table 2 illustrates ADM-W++ achieves markedly superior performance compared to ADM, ADM-IP, and ADM-ES, regardless of datasets or timesteps. It is worth noting that in the 20-step and 30-step sampling tasks on CIFAR-10, ADM-W++ attains the remarkable FID, which is less than One half of ADM’s. Specifically, “-” in tables indicates that the original paper of the comparison models did not report results for the task.
To prove the superiority of W++, we select the improved models for exposure bias, TS-DPM (Li et al., 2024) and MDSS (Ren et al., 2024) as the comparison models, with DDPM (Ho et al., 2020) and IDDPM (Nichol and Dhariwal, 2021) as the baseline models. Table 3 clearly shows that regardless of the timestep or baseline, W++ achieves much higher generation quality than TS-DPM and MDSS. In particular, in the comparative experiments with TS-DPM, the performance of W++ is far superior to that of TS-DPM, which further highlights the superiority of W++.
To demonstrate the universality of W++, we selected higher-resolution datasets for testing, namely ImageNet $\times 128$ 128 and LSUN bedroom $\times 256$ 256. Meanwhile, we selected ADM and IDDPM as baseline models for these two datasets respectively, to align as closely as possible with the experimental settings of the comparison model described in its original paper. Table 4 makes it clear that, regardless of the dataset or the timestep, W++ achieves better metrics than both the baseline model and the comparison model, which further highlights the generality of W++.
Table 4. FID on ImageNet $\times 128$ 128 and LSUN Bedroom $\times 256$ 256 using ADM and IDDPM.
ADM & ImageNet IDDPM & Bedroom
Method 20 30 50 20 30 50
baseline 12.48 8.04 5.15 18.63 12.99 8.50
baseline-ES 9.86 6.35 4.33 10.36 6.69 4.70
baseline-W++ 8.81 5.88 4.22 9.37 6.24 4.66
Table 5. FID on CIFAR-10 using A-DPM and EA-DPM.
CIFAR10 (LS) CIFAR10 (CS)
$T^{\prime}$ 10 25 50 10 25 50
DDPM, $\tilde{\beta}_{n}$ 44.45 21.83 15.21 34.76 16.18 11.11
DDPM, $\beta_{n}$ 233.41 125.05 66.28 205.31 84.71 37.35
A-DPM 34.26 11.60 7.25 22.94 8.50 5.50
A-DPM-W++ 12.38 6.63 4.52 11.61 4.40 3.62
NPR-DPM 32.35 10.55 6.18 19.94 7.99 5.31
LA-NPR-DPM 25.59 8.84 5.28 - - -
NPR-DPM-W++ 10.86 5.76 4.11 10.18 4.07 3.44
SN-DPM 24.06 6.91 4.63 16.33 6.05 4.17
LA-SN-DPM 19.75 5.92 4.31 - - -
SN-DPM-W++ 11.73 4.73 3.78 12.53 4.51 3.47
5.2. Results on A-DPM and EA-DPM
To further evaluate the versatility of W++, we choose the improved diffusion models A-DPM (Analytic-DPM) (Bao et al., 2022b) and EA-DPM (Extended-Analytic-DPM) (Bao et al., 2022a). The former deduced the analytic form of the optimal reverse variance and the latter used neural networks to estimate the optimal covariance of the conditional Gaussian distribution in the inverse process. Specifically, we tested two advanced models in EA-DPM: NPR-DPM and SN-DPM, where the “NPR” and “SN” symbols represent two different methods for estimating optimal variance under different conditions in EA-DPM. Meanwhile, we conducted various training strategies, employing the linear schedule of $\beta_{n}$ (LS) (Ho et al., 2020), cosine schedule of $\beta_{n}$ (CS) (Nichol and Dhariwal, 2021) respectively. Finally, we chose LA-DDPM (Zhang et al., 2023) as comparison model because this work made improvements on EA-DPM.
Table 5 shows that W++ outperforms all baseline and comparison models by achieving lower FID scores, regardless of the choices of estimation method, training strategy, or sampling timestep, which not only shows W++ can significantly improve generation quality, but also that W++ is more advanced than other improvements. Similar to the pattern observed with ADM, W++ shows the same degree of performance on any timestep and dataset. These results further confirm the effectiveness and versatility of W++.
5.3. Results on DDIM, EDM, FPGM++, and AMED
To further verify the effectiveness and versatility of W++, we select several deterministic sampling models DDIM (Song et al., 2021a), EDM (Karras et al., 2022), FPGM++ (Xu et al., 2023), and AMED (Zhou et al., 2024). We choose Analytic-DPM (Bao et al., 2022b) as the backbone model for DDIM and the Euler method as the solver for EDM and PFGM++. Unlike the noise prediction network, EDM’s neural network restores the original data, PFGM++ is an improved model of the Poisson Flow Generative Model (Xu et al., 2022), and AMED significantly mitigates truncation errors using the median theorem, all of which are crucial for testing the generality of W++. In particular, for DDIM, we use the usual sampling time step as $T^{\prime}$ as in previous experiments, but for EDM, PFGM ++ and AMED, we choose Neural Function Evaluations (NFE) (Vahdat et al., 2021) as $T^{\prime}$.
Table 6 clearly shows that W++ can significantly improve the generation quality of DDIM; For AMED using the mean value theorem and distillation techniques, W++ can still reduce FID by mitigating exposure bias. Although the EDM illustrates the design space of the diffusion models and PFGM++ is a Poisson Flow Generation Model, the results in Table 7 also maintain the same trend, with W++ showing better generation performance than the baseline model and the comparison model. The above experiments show for different deterministic sampling methods, different time steps, and different neural network types, W++ can always significantly improve the generation quality of baseline models, which further proves the effectiveness and versatility of W++.
Particularly, in all the above experiments, $t_{\rm mid}$ is selected as 0.2.
Table 6. FID on CIFAR-10 using DDIM and AMED.
DDIM AMED
$T^{\prime}$ 10 25 50 5 7 9
Baseline 26.43 9.96 6.02 7.59 4.36 3.67
Baseline-ES 22.64 7.24 4.58 7.53 4.36 3.66
Baseline-W++ 19.50 6.50 4.46 7.01 4.22 3.54
Table 7. FID on CIFAR-10 using EDM and PFGM++.
EDM PFGM++
$T^{\prime}$ 13 21 35 13 21 35
Baseline 10.66 5.91 3.74 12.92 6.53 3.88
Baseline-ES 6.59 3.74 2.59 8.79 4.54 2.91
Baseline-W++ 4.68 2.84 2.13 6.62 3.67 2.53
Table 8. Ablation study of the two regulation schemes.
[…]
ADM 10.55 8.96 0.65 0.53
ADM-low 7.49 9.24 0.66 0.57
ADM-high 7.58 9.67 0.66 0.57
ADM-W++ 4.62 9.74 0.67 0.60
5.4. Qualitative comparison
To visually show the impact of W++ on the generation quality of DPMs, we set the same random number seed and sampling steps for ADM and ADM-W++ during sampling to ensure that the sampling trajectories of two models are as similar as possible. Fig. 7 clearly shows that due to exposure bias, the generated samples by ADM exhibit unnatural phenomena such as over-smoothing, over-exposure, or being too dark. However, ADM-W++ significantly alleviates these problems, resulting in more natural-looking generated images. Specifically, ADM-W++ combats the reduction of frequency-domain energy caused by exposure bias by dynamically compensating the low-frequency and high-frequency energy of the image, which further proves the superiority of W++.
Figure 7. Qualitative comparison between ADM (first row) and ADM-W++ (second row) using 20 steps.
Table 9. Time for a batch generation on LSUN bedroom $\times 256$ 256 using one GeForce RTX 3090.
Type IDDPM IDDPM-W++ One W++ denoise One w++
Time 17.35 17.42 0.873 0.003
5.5. Ablation study
Synergistic Effect. To explore the synergy of W++, we separately adjusted the low and high-frequency components, denoted as ADM-low and ADM-high, respectively, and then adjusted both components simultaneously, denoted as ADM-W++. For a comprehensive evaluation of fidelity, diversity, and accuracy, we selected Inception Score (IS), recall, and precision as additional metrics. As shown in Table 8, adjusting any single frequency component—either low or high—substantially enhances the generation quality. However, the joint adjustment of both low and high-frequency components consistently outperforms individual adjustments, highlighting the synergistic effect of combining both frequency subband.
[…]
Hyperparameter Insensitivity. To demonstrate the insensitivity of W++ to hyperparameters, we present the detailed process of parameter search. Initially, we adjusted only the low-frequency component of baselines to find the optimal parameter $w_{l}^{*}$. Then, based on the optimal $w_{l}^{*}$, we fine-tuned the high-frequency components to determine the optimal $w_{h}^{*}$. Fig. 8 clearly shows that W++ achieves consistent performance improvements over a wide range of $w_{l}$ and $w_{h}$, which indicates the method’s insensitivity to hyperparameters. Notably, Fig. 8 also shows the FID curves always follow the pattern of first decreasing and then increasing, and the inflection point corresponds to the optimal parameter. This explicit characteristic of the minimum point enables us to quickly lock in the optimal parameter, further verifying the practicality of W++.
(a) Search experiments of $w_{l}$
(b) Search experiments of $w_{h}$
Figure 8. Hyperparameters search experiments on CIFAR-10 using A-DPM and EA-DPM with 25 steps.
6. Conclusion
In this work, we analyze and mitigate exposure bias of diffusion probabilistic models in the wavelet domain for the first time. We find that the subband energy of predicted samples in the reverse process follows distinct reduction patterns, thus we propose a simple yet effective frequency regulation mechanism that dynamically enhances the energy of low and high-frequency subbands during sampling. We also find that the subband energy of reconstructed samples is consistently lower than that of original data samples, which strongly supports our theoretical analysis. Based on this, we further derive the analytical form of exposure bias. Notably, our method is training-free, plug-and-play, and significantly improves the generation quality of various diffusion probabilistic models and frameworks, with negligible computational cost. Moreover, our method was extensively tested on datasets of varying resolutions, different samplers, and diverse time steps. Superior results demonstrate its significant outperform over other exposure bias-improved models, highlighting its advancement.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under the Grant No. 62176108 and Supercomputing Center of Lanzhou University.
Appendix A
In this section, we give a detailed proof of the Eq. (13). For ease of understanding, we will rewrite the necessary formulas here and we state that all the ${\bm{\epsilon}}$ proposed in this section follow the standard Gaussian distribution. We know that the perturbed sample in the forward process is
(20) ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}.$
Assumption 111 is expressed as
(21) ${\bm{x}}^{0}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)=\gamma_{t}{\bm{x}}_{0}+\phi_{t}{\bm{\epsilon}}_{t}.$
The predicted sample in the reverse process is
(22) $\hat{{\bm{x}}}_{t-1}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t}+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1},$
where ${\bm{x}}^{0}_{\bm{\theta}}({\bm{x}}_{t},t)=\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{\epsilon}}_{\bm{\theta}}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}$. We can substitute Eq. (20) and Eq. (21) into Eq. (22) to obtain
(23) $\hat{{\bm{x}}}_{t-1}$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}(\gamma_{t}{\bm{x}}_{0}+\phi_{t}{\bm{\epsilon}}_{t})+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}(\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}})$
$+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1}$
$=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}}{1-\bar{\alpha}_{t}}{\bm{x}}_{0}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\phi_{t}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{t}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}{\bm{x}}_{0}$
$+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{1-\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1}$
$=\left(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}}{1-\bar{\alpha}_{t}}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}\right){\bm{x}}_{0}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\phi_{t}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{t}$
$+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{1-\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+\sqrt{\tilde{\beta}_{t}}{\bm{\epsilon}}_{1}.$
Now, we deal with the first part of Eq. (23):
(24) $\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}}{1-\bar{\alpha}_{t}}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}\gamma_{t}+\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}$
$=\frac{\sqrt{\bar{\alpha}_{t-1}}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t}}$
$=\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}.$
Then, we need to construct an auxiliary term. Therefore, we discard $\gamma_{t}$ in Eq. (24):
(25) $\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}1-\alpha_{t}\bar{\alpha}_{t-1}\big{)}}{1-\bar{\alpha}_{t}}=\sqrt{\bar{\alpha}_{t-1}}.$
Since $1-\alpha_{t}>0,\gamma_{t}\leqslant 1$, according to Eq. (24) and Eq. (25), we can naturally obtain
$\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}\leqslant\sqrt{\bar{\alpha}_{t-1}}$
We can definitely define a new coefficient $\gamma_{t-1}\leqslant 1$ such that
(26) $\frac{\sqrt{\bar{\alpha}_{t-1}}\big{(}(1-{\alpha}_{t})\gamma_{t}+{\alpha_{t}}(1-\bar{\alpha}_{t-1})\big{)}}{1-\bar{\alpha}_{t}}=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}.$
For the second part of Eq. (23), we can always split it into
(27) $Var(\hat{{\bm{x}}}_{t-1})$ $=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\sqrt{1-\bar{\alpha}_{t-1}})^{2}+\tilde{\beta}_{t}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+(\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\sqrt{1-\bar{\alpha}_{t-1}})^{2}$
$+\frac{(1-\bar{\alpha}_{t-1})(1-\alpha_{t})}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{\alpha_{t}(1-\bar{\alpha}_{t-1})^{2}}{1-\bar{\alpha}_{t}}+\frac{(1-\bar{\alpha}_{t-1})(1-\alpha_{t})}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{\alpha_{t}(1-\bar{\alpha}_{t-1})^{2}+(1-\bar{\alpha}_{t-1})(1-\alpha_{t})}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{(1-\bar{\alpha}_{t-1})[\alpha_{t}(1-\bar{\alpha}_{t-1})+(1-\alpha_{t})]}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{(1-\bar{\alpha}_{t-1})[\alpha_{t}-\bar{\alpha}_{t}+1-\alpha_{t}]}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+\frac{(1-\bar{\alpha}_{t-1})[1-\bar{\alpha}_{t}]}{1-\bar{\alpha}_{t}}$
$=(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}+1-\bar{\alpha}_{t-1}.$
Based on Eqs. (26) and  (27), we can obtain
(28) $\hat{{\bm{x}}}_{t-1}=\gamma_{t-1}\sqrt{\bar{\alpha}_{t-1}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t-1}+(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\phi_{t})^{2}}{\bm{\epsilon}}_{t-1}$
Appendix B
In this section, we present the parameter settings for the important experiments in this paper, as shown in Tables 10, 11, 12, 13, and 14.
Table 10. Settings on CIFAR-10 and Image-Net using ADM.
$w_{l}/w_{h}$ CIFAR-10 Image-Net
$T^{\prime}$ 20 30 50 20 30 50
ADM $w_{l}$ 1.013 1.008 1.0036 0.050 0.040 0.028
ADM $w_{h}$ 1.064 1.034 1.015 0.997 0.998 1.001
Table 11. Settings on CIFAR-10 using DDPM and IDDPM.
$w_{l}/w_{h}$ DDPM IDDPM
$T^{\prime}$ 10 20 30 100
Baseline $w_{l}$ 1.068 1.019 1.009 1.0011
Baseline $w_{h}$ 1.250 1.140 1.042 1.0060
Table 12. Settings on CIFAR-10 using A-DPM, EA-DPM.
$w_{l}/w_{h}$ CIFAR10 (LS) CIFAR10 (CS)
$T^{\prime}$ 10 25 50 10 25 50
A-DPM $w_{l}$ 0.132 0.049 0.03 0.072 0.032 0.008
A-DPM $w_{h}$ 1.11 1.038 1.019 1.202 1.046 1.018
NPR-DPM $w_{l}$ 0.132 0.048 0.024 0.066 0.03 0.007
NPR-DPM $w_{h}$ 1.105 1.034 1.015 1.192 1.045 1.017
SN-DPM $w_{l}$ 0.109 0.043 0.025 0.052 0.027 0.009
SN-DPM $w_{h}$ 1.013 1.005 1.005 1.101 1.019 1.01
Table 13. Settings on CIFAR-10 using DDIM and AMED.
$w_{l}/w_{h}$ DDIM AMED
$T^{\prime}$ 10 25 50 5 7 9
Baseline $w_{l}$ 0.0 0.0 0.0 0.0014 0.0012 0.0027
Baseline $w_{h}$ 1.21 1.037 1.011 0.9955 0.9932 0.9985
Table 14. Settings on CIFAR-10 using EDM and PFGM++.
$w_{l}/w_{h}$ EDM PFGM++
$T^{\prime}$ 13 21 35 13 21 35
Baseline $w_{l}$ 0.036 0.016 0.007 0.037 0.016 0.006
Baseline $w_{h}$ 1.087 1.054 1.022 1.095 1.057 1.025
Table 15. The search experiment of $w_{l}$ on CIFAR-10(LS) using A-DPM with 25 steps.
$w_{l}$ 0.0 0.03 0.04 0.049 0.05 0.06 0.07
Time 11.60 9.00 8.59 8.46 8.47 8.59 8.99
Appendix C
We emphasize that W++ can quickly search for the optimal parameters, and in this section, we provide more favorable evidence. Specifically, $w_{t}^{l}$ and $w_{t}^{h}$ are adaptively determined by the inverse variance, and and regulated by $w_{l}$, $w_{h}$, and $t_{mid}$, which are searched efficiently. In particular, we select CIFAR-10 as the test dataset and A-DPM (Bao et al., 2022b) as the baseline model, and then the 25-step sampling task will be performed.
(1) $w_{l}$ and $w_{h}$. Due to the method’s insensitivity to parameters, the parameter search process is fast via the two-stage search. Firstly, a coarse search with a step size of 0.010.010.01 was performed. After finding a turning point in FID near 0.050.050.05, a fine search with a step size of 0.0010.0010.001 was conducted, quickly determining the optimal value as 0.049, as shown in Table 15.
[…]
(3) $\sigma_{t}$. $\sigma_{t}$ is known and directly set to the empirical inverse variance of the baseline model.